Circular No. 7976

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V4743 SAGITTARII = ANOTHER NOVA IN SAGITTARIUS N. N. Samus, Institute of Astronomy, Moscow, informs us that the

designation V4743 Sgr has been given to this nova (cf. IAUC 7975).

T. Kato, Kyoto University, reports that six prediscovery Tri-X 400 films taken by T. Hishikura (Yokohama, Kanagawa, Japan) around Sept. 18.465 UT with a 200-mm f/2.8 camera lens shows the nova at mag 5.5. CCD photometry by J. D. West, Mulvane, KS: Sept. 23.070, V = 6.37. Further visual magnitude estimates: Sept. 21.448, 5.5 (A. Beresford, Adelaide, S. Australia); 21.993, 6.0 (J. Garcia, Rama Caida, Argentina); 22.10, 6.1 (A. Hale, Cloudcroft, NM; moonlight); 22.354, 6.2 (J. Bedient, Honolulu, HI); 22.523, 6.2 (A. Pearce, Nedlands, W. Australia); 22.912, 6.3 (A. Amorim, Florianopolis, Brazil; moonlight); 23.322, 6.3 (Bedient); 23.365, 5.8 (S. J. O’Meara, Volcano, HI); 23.496, 6.1 (Pearce); 23.867, 6.2 (L. Parmeggiani, Bray-et-Lu, France).

SUPERNOVA 2002fk IN NGC 1309 K. Ayani, Bisei Astronomical Observatory (BAO); and H. Yamaoka,

Kyushu University, write: “A low-resolution spectrum (range 500–700 nm) of SN 2002fk (cf. IAUC 7973), obtained with BAO 1.01-m telescope on Sept. 20.7 UT, indicates that it is a type-Ia supernova before maximum light. Adopting the NASA/IPAC Extragalactic Database recession velocity of 2135 km/s, the expansion velocity deduced from Si II (rest 635.5 nm) is ∼ 10900 km/s.”

CCD photometry, unfiltered unless otherwise noted: Sept. 12.110 UT, [18.0 (L. A. G. Monard, Bronberg Ridge, near Pretoria, S. Africa); 19.848, 14.5 (Monard); 20.095, 14.4 (Monard); 20.421, V = 14.1 (J. D. West, Mulvane, KS); 21.003, 14.1 (Monard); 21.998, 13.9 (Monard); 23.099, 13.7 (Monard).

V838 MONOCEROTIS A. M. Watson and R. Costero, Instituto de Astronomía, Universidad

Nacional Autónoma de México, report the following magnitudes obtained at the Observatorio Astronómico Nacional 2.1-m telescope (+ CID camera) on Sept. 19 UT: [7.7 μm] = 2.1:, [8.7 μm] = 1.5, [9.7 μm] = 1.3, [10.3 μm] = 1.3, [11.7 μm] = 1.2, [12.5 μm] ∼ 1.3; [18 μm] = [1.0 (not detected). The star has brightened in the range 8.7–11.9 μm by ∼ 3 mag since January (cf. IAUC 7831).

2002 September 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green